OpenStudy (anonymous):

what is the simplified form of 4/y^2-5/3y?

OpenStudy (stamp):

To simplify the expression, you need to combine the two terms. To combine the two terms, each term needs a common denominator.

OpenStudy (stamp):

\[4/y^2-5/3y\]\[12/3y^2-5y^2/3y^2\]simplified form = http://www.wolframalpha.com/input/?i=simplify+4%2Fy^2-5%2F3y

OpenStudy (anonymous):

=5y-12/3y^2 =4y-15/3y^2 =12-5y/3y^2 =15-4y/3y^2

OpenStudy (anonymous):

Which one is it?

OpenStudy (anonymous):

HEEEEELLLLP

OpenStudy (stamp):

Look at the attached image. This is a screen capture of the link provided earlier.